OpenStudy (anonymous):

What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12?

OpenStudy (anonymous):

I have no idea what to do..

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Come on walk me through it I need it now!

OpenStudy (anonymous):

d=-6, a11=12 \[a11=a1+(11-1)d=a1+10(-6)\]

OpenStudy (anonymous):

a1-60=12 find a1 \[a43=a1+42d=a1+42(-6)\]